# Instruction Matters, a Simple yet Effective Task Selection Approach in Instruction Tuning for Specific Tasks

### v1

@@ Abstract
Instruction tuning has shown its ability to not only enhance zero-shot generalization across various tasks but also its effectiveness in improving the performance of specific tasks. A crucial aspect in instruction tuning for a particular task is a strategic selection of related tasks that offer meaningful supervision, thereby enhancing efficiency and preventing performance degradation from irrelevant tasks. Our research reveals that leveraging instruction information alone enables the identification of pertinent tasks for instruction tuning. This approach is notably simpler compared to traditional methods that necessitate complex measurements of pairwise transferability between tasks or the creation of data samples for the target task. Furthermore, by additionally learning the unique instructional template style of the meta-dataset, we observe an improvement in task selection accuracy, which contributes to enhanced overall performance. Experimental results demonstrate that training on a small set of tasks, chosen solely based on the instructions, leads to substantial performance improvements on benchmarks like P3, Big-Bench, NIV2, and Big-Bench Hard. Significantly, these improvements exceed those achieved by prior task selection methods, highlighting the efficacy of our approach.111Our code and dataset will be released after the review process.
Instruction Matters, a Simple yet Effective Task Selection Approach in
Instruction Tuning for Specific Tasks
@@ 1 Introduction
Recently, instruction tuning has gained attention as an innovative approach for improving zero-shot performance Sanh et al. (2022); Wei et al. (2022); Wang et al. (2022); Ouyang et al. (2022); Taori et al. (2023); Chiang et al. (2023). This method offers the advantage of improving a model’s generalization capabilities to unseen tasks by training on diverse tasks accompanied by instructions. The robustness of instruction tuning improves as the diversity of training tasks increases Sanh et al. (2022); Wang et al. (2022); Longpre et al. (2023); Chung et al. (2022). In this light, recent studies have concentrated on broadening the diversity and increasing the number of tasks within the meta-dataset Triantafillou et al. (2020) used for instruction tuning. For example, Wang et al. (2022) introduce a collection of over 1,600 tasks, and Chung et al. (2022) expand to 1,836 tasks through the integration of additional datasets. Moreover, Wang et al. (2023c); Yin et al. (2023); Lee et al. (2023) automatically generate a substantial volume of instruction tuning data using Large Language Models (LLMs) instead of human creation.
@@
Besides applying instruction tuning for general unseen tasks, there is also growing interest in instruction tuning as a methodology to improve the performance of specific unseen tasks Wang et al. (2023a); Zhang et al. (2023); Feng et al. (2023). Instruction tuning focusing on specific tasks trains not all but only informative tasks with instruction format. This strategy is based on the insight that not all tasks are helpful to specific tasks, and some tasks could even lead to performance degradation due to negative transfer during multi-task training Muennighoff et al. (2023); Wang et al. (2023b); Zhou et al. (2023); Jang et al. (2023); Kim et al. (2023a); Shi et al. (2023). However, selecting relevant tasks for training presents a significant challenge. First, manually reviewing through the vast array of datasets for instruction tuning is not feasible. Additionally, discerning the relevance of certain tasks in the training dataset to the target task is often ambiguous.
@@
To address this challenge, studies have been conducted to automatically quantify the relevance between tasks Lin et al. (2022); Ye et al. (2022); Jang et al. (2023); Paranjape et al. (2023); Kim et al. (2023a). These methods are primarily divided into two categories. The first is assessing pairwise task transferability by training models on one task and then evaluating their performance on another Vu et al. (2020a); Poth et al. (2021a); Zhou et al. (2023); Kim et al. (2023a). The second is calculating similarity scores between tasks using a small sample of the test dataset Lin et al. (2022); Ye et al. (2022); Jang et al. (2023); Paranjape et al. (2023). However, the former approach, which measures pairwise transfer for every task, can be exceedingly time-consuming and compute-intensive. Moreover, the latter approach has limitations in that it necessitates constructing data (instances) for unseen target tasks, contradicting the zero-shot setting condition in instruction tuning.
@@
In this study, we explore a straightforward yet effective method for choosing relevant tasks for instruction tuning. We focus on the feature of instruction tuning, where the instructions222Following Sanh et al. (2022); Wei et al. (2022); Wang et al. (2022), instruction means prompt, template, and task description without instance. define the characteristics of each task. In this vein, we introduce an Instruction-based Task selector (InsTa) that leverages instruction similarity scores to determine task relevance and select informative tasks for training. Through experiments, we discover that this simplified instruction-only selection method adequately identifies related tasks and significantly improves performance. Moreover, by further aligning InsTa through training on the instruction style and format of a meta-dataset, it is able to closely understand the instructional nuances, achieving significant performance improvements.
@@
A notable finding from our study is that task selection based exclusively on instruction surpasses previous methods Lin et al. (2022); Ye et al. (2022); Jang et al. (2023); Paranjape et al. (2023) that incorporate both instructions and instances. Moreover, instruction information alone shows a high correlation with task similarity as determined by complex pairwise transfer methods Zhou et al. (2023) and even achieves slightly better average zero-shot performance. This indicates that the instruction-based approach for selecting related tasks is not only easily applicable but also highly effective.
@@
We introduce InsTa which simplifies task selection for instruction tuning by using only instruction similarity scores, obviating the need for instances.
By aligning InsTa with the instruction styles and formats of meta-datasets, we significantly improve performance, demonstrating the importance of understanding instructional nuances.
Our study reveals that using only instructions not only surpasses prior methods that employing both instructions and instances but also parallels the trends of pairwise transfer methods without their associated complexities, demonstrating our method’s effectiveness.
@@ 2.1 Instruction Tuning: Generalist or Specialist
According to Kung and Peng (2023), instruction tuning as a generalist can be categorized based on its objectives: generalizing to "unseen tasks" and generalizing to "unseen instructions". Early studies propose the approach of generalizing to unseen tasks, which involves training on various natural language processing (NLP) tasks and evaluating on unseen tasks Wei et al. (2022); Sanh et al. (2022); Wang et al. (2022). However, task-based instruction tuning presents the difficulty of defining every task that humans use. In response, a methodology that trains LLMs on diverse instructions without clear distinctions for specific tasks has been proposed, aiming to generalize to unseen instructions, rather than unseen tasks Ouyang et al. (2022); Taori et al. (2023); Chiang et al. (2023).
@@ 2.2 Quantifying Task Relationship
The research on identifying task relationships has been extended to the field of instruction tuning as well. Zhou et al. (2023); Kim et al. (2023a) measure pairwise task transfer between every task pair to identify helpful source tasks on specific target tasks. On the other hand, Jang et al. (2023); Paranjape et al. (2023) have calculated task similarity using only a few data samples from training and evaluation tasks instead of training and evaluating every task individually.
Previous task selection approach, which automatically measure task similarities, are either time-consuming due to the need to train and evaluate every task pair or still rely on having data samples for the target tasks, even in reduced quantities. Our method simplifies this process by evaluating task relationships solely through instructions, eliminating the need for laborious measurements across every task or generating test data, provided the tasks are adequately described through instructions. For more detailed descriptions of our method’s practicalities compare to prior works, please refer to Appendix A.
@@ 3.1 Formulation
where denotes the instruction of the target task , and denotes the instruction of some arbitrary task chosen for similarity assessment. For measuring similarity, we employ cosine similarity, and for the embedding function , we utilize the Sentence Transformer Reimers and Gurevych (2019) as an off-the-shelf embedding model, following Jang et al. (2023). For more specific details, please refer to the Appendix B.
@@ 3.2 Aligning Instruction-based Task Selector with Meta-dataset
Off-the-shelf embedding model often lacks the capability to accurately identify related tasks through instructions, since they are not trained on the unique instruction styles present in meta-datasets. To mitigate this issue, our approach includes an additional aligning process that fine-tunes our selector model on the distinctive instruction styles of each meta-dataset. For training, we select a random instruction from a given task as a positive sample and designate instructions from different task clusters as negative samples.333The rationale behind not utilizing task instructions of the same task cluster as positive or negative samples is twofold: in the former case, even though tasks may have the same task type, they can be differentiated strictly, and in the latter case, false negatives may occur, even within same task type. The training objective is as follows:
@@ 4.1 Dataset
We conduct experiments on two representative instruction tuning meta-datasets: P3 (Public Pool of Prompts) Sanh et al. (2022); Bach et al. (2022) and NIV2 (SuperNaturalInstructions V2) Wang et al. (2022). P3 is a meta-dataset comprised of 12 task clusters. It contains a total of 35 tasks across 8 task clusters for training and 11 tasks across 4 task clusters for held-out evaluation. Note that each task in P3 includes 11.7 instructions on average. Conversely, NIV2 encompasses 72 task clusters of English tasks. The training set comprises a total of 756 tasks across 60 clusters, while the held-out evaluation tasks include 119 tasks across 12 task clusters. In the case of NIV2, each task consists of a single instruction. We additionally evaluate BIG-Bench bench authors (2023) and BIG-Bench Hard (BBH) Suzgun et al. (2023) as supplementary evaluation datasets. For more detailed information, please refer to the Table 1 and Appendix D.1.
@@ 4.2 Task Selector Setup
instructions feature human-crafted, human-readable Task Definition, Positive Task Examples, Negative Task Examples, and Explanation. We utilize the Task Definition for training the task selector. Moreover, since NIV2 has only one instruction per task, there are no positive samples for training the task selector. To address this, we generate instructions for all NIV2 tasks using GPT-4 OpenAI (2023). Specifically, we generate a paraphrased one and employ it as a positive sample for training the task selector. For more information about a query used for GPT-4 API calls and generated instructions, please refer to the Appendix E.1.
@@ 4.3 Training Setup
Following previous studies Sanh et al. (2022); Wang et al. (2022), we employ all tasks except for those held out for evaluation as our training set. Unlike traditional approaches that train on all tasks, our strategy specifically trains only the top- tasks deemed most informative for each target task.
@@
We use the T5 LM-adapted model(3B) Lester et al. (2021) as our base model and train for 3 epochs with a constant learning rate of 1e-4 and 5e-5 for P3 and NIV2, respectively. We configure validation from training datasets and select the model that shows the best performance in the validation. Our experiment is in a true zero-shot setting, where no samples from the held-out task are used for checkpoint selection. Please refer to the Appendix C for more detail.
@@
For the P3 dataset, we use the following baselines. GPT-3(175B) is an autoregressive LM that has shown remarkable ability in following demonstrations provided in its instructions Brown et al. (2020). T0-11B / T0-3B have same architecture as T5 but trained on millions of samples from 35 different P3 tasks Sanh et al. (2022). T5(3B) is text-to-text pretrained LM without instruction tuning Raffel et al. (2020). T5(3B) + Random is trained on 5 random tasks from P3. T5(3B) + Pairwise Transfer is trained on top-5 tasks with the highest transferability scores demonstrated in Zhou et al. (2023), representing pairwise transfer task selection. T5(3B) + PE w/RoE denotes a model that selects relevant task by measuring cosine similarity between training and evaluation data samples Jang et al. (2023).
@@ 4.4 Evaluation Setup
For P3 evaluation, following Sanh et al. (2022), we apply rank classification and measure the performance of the model for each target instruction. Then, we calculate the average performance of instructions per task. Note that each target task in P3 has 10.09 instructions on average. For NIV2 evaluation, we follow the same evaluation protocol as in Wang et al. (2022) and report ROUGE-L Lin (2004) score. We adopt greedy decoding with a maximum generation length of 256.
@@ 5.1 P3 Results
Table 2 shows experimental results on the 11 unseen datasets from P3. Compared to T0-3B, which is fully instruction-tuned on all 35 tasks from P3, the T5(3B) + Random model shows inferior performance, but both T5(3B) + InsTa model and T5(3B) + InsTa model outperform 10 tasks out of 11 tasks, with each exhibiting a performance gap of 4.83% and 7.10% on average, respectively. They also show superior performance compared to T5(3B) + PE w/ RoE, highlighting the instruction itself is sufficient enough to choose informative tasks without the use of data samples. Moreover, T5(3B) + InsTa outperforms T5(3B) + InsTa across 10 tasks, suggesting further aligning task selector on instructions from the meta-dataset increases the precision of task selection. Additionally, our approach exhibits marginally superior performance compared to T5(3B) + Pairwise Transfer, suggesting that instruction-only task selection can effectively identify related tasks without the need for exhaustive pairwise evaluations. Notably, we observe a strong correlation between pairwise task transferability and our task selection approach, further elaborated in Appendix F.
@@ 5.2 Big-Bench Results
Table 3 evaluates the performance of instruction-based task selection on the 14 Big-Bench tasks employing P3 held-in tasks in Section 5.1 as the training dataset. We compare our method to baselines T0-3B and T5(3B) + Cos PE of Jang et al. (2023). Our model, T5(3B) + InsTa, surpasses T0-3B in 12 out of 14 tasks, achieving an average performance increase of 5.94%. Tasks like Movie D. and Syllogism, where results are comparable to random guesses, indicate informative tasks are insufficient in P3. When compared to Cos PE Jang et al. (2023), our T5(3B) + InsTa model shows improvements in most tasks, with an average increase of 2.07%. Notably, it also surpasses T0-11B in a majority of tasks. These findings demonstrate that instruction-based task selection enables effective training even with a small number of relevant tasks.

### v2

# Instruction Matters: A Simple yet Effective Task Selection for Optimized Instruction Tuning of Specific Tasks

@@ Abstract
Instruction tuning has been proven effective in enhancing zero-shot generalization across various tasks and in improving the performance of specific tasks. For task-specific improvements, strategically selecting and training on related tasks that provide meaningful supervision is crucial, as this approach enhances efficiency and prevents performance degradation from learning irrelevant tasks. In this light, we introduce a simple yet effective task selection method that leverages instruction information alone to identify relevant tasks, optimizing instruction tuning for specific tasks. Our method is significantly more efficient than traditional approaches, which require complex measurements of pairwise transferability between tasks or the creation of data samples for the target task. Additionally, by aligning the model with the unique instructional template style of the meta-dataset, we enhance its ability to granularly discern relevant tasks, leading to improved overall performance. Experimental results demonstrate that training on a small set of tasks, chosen solely based on the instructions, results in substantial improvements in performance on benchmarks such as P3, Big-Bench, NIV2, and Big-Bench Hard. Significantly, these improvements surpass those achieved by prior task selection methods, highlighting the superiority of our approach.111Code, model checkpoints, and data resources are available at https://github.com/CHLee0801/INSTA.
Instruction Matters: A Simple yet Effective Task Selection for
Optimized Instruction Tuning of Specific Tasks
Changho Lee1††thanks: Equal contribution. Janghoon Han1 Seonghyeon Ye2 Stanley Jungkyu Choi1 Honglak Lee1 Kyunghoon Bae1 1LG AI Research 2KAIST {changho.lee,janghoon.han}@lgresearch.ai
@@ 1 Introduction
Recently, instruction tuning has gained attention as an innovative approach for improving zero-shot performance Sanh et al. (2022); Wei et al. (2022); Wang et al. (2022); Ouyang et al. (2022); Taori et al. (2023); Chiang et al. (2023). This method offers the advantage of improving a model’s generalization capabilities to unseen tasks by training on diverse tasks accompanied by instructions. The robustness of instruction tuning improves as the diversity of training tasks increases Sanh et al. (2022); Wang et al. (2022); Longpre et al. (2023); Chung et al. (2022). In this light, recent studies have concentrated on broadening the diversity and increasing the number of tasks within the meta-dataset Triantafillou et al. (2020) used for instruction tuning Wang et al. (2022); Chung et al. (2022); Wang et al. (2023c); Yin et al. (2023); Lee et al. (2023).
@@
Besides applying instruction tuning for general unseen tasks, there is also growing interest in instruction tuning as a methodology to improve the performance of specific unseen tasks Wang et al. (2023a); Zhang et al. (2023); Feng et al. (2023). Instruction tuning focuses on specific tasks and trains not all but only informative tasks with instruction format. This strategy is based on the insight that not all tasks are helpful to specific tasks, and some tasks could even lead to performance degradation due to negative transfer during multi-task training Muennighoff et al. (2023); Wang et al. (2023b); Zhou et al. (2023); Jang et al. (2023); Kim et al. (2023a); Shi et al. (2023). However, selecting relevant tasks for training presents a significant challenge. First, manually reviewing through the vast array of datasets for instruction tuning is not feasible. Additionally, discerning the relevance of certain tasks in the training dataset to the target task is often ambiguous.
@@
To address this challenge, studies have been conducted to automatically quantify the relevance between tasks Lin et al. (2022); Ye et al. (2022); Jang et al. (2023); Paranjape et al. (2023); Kim et al. (2023a). These methods are primarily divided into two categories. The first assesses pairwise task transferability by training models on one task and evaluating their performance on another Vu et al. (2020a); Poth et al. (2021a); Zhou et al. (2023); Kim et al. (2023a). The second calculates similarity scores between tasks by comparing small samples from each task’s dataset Lin et al. (2022); Ye et al. (2022); Jang et al. (2023); Paranjape et al. (2023). However, the former approach, which measures pairwise transfer for every task, can be exceedingly time-consuming and compute-intensive. Moreover, the latter approach still necessitates the construction of data for unseen target tasks, which not only introduces additional burdens but also does not fully align with the fundamental goal of instruction tuning, which is to enhance zero-shot capabilities.
@@
In this study, we explore a straightforward yet effective method for choosing relevant tasks for optimized instruction tuning. We focus on the feature of instruction tuning, where the instructions222Following Sanh et al. (2022); Wei et al. (2022); Wang et al. (2022), instruction means prompt, template, and task description without instance. define the characteristics of each task. Building upon this, we introduce an Instruction-based Task selector (InsTa) that leverages instruction similarity scores to determine task relevance and select informative tasks for training. Through experiments, we discover that this simplified instruction-only selection method adequately identifies related tasks and significantly improves performance. Moreover, by further aligning InsTa through training on the instruction style and format of a meta-dataset, it is able to closely understand the instructional nuances, achieving significant performance improvements.
@@
A notable finding from our study is that task selection based exclusively on instructions surpasses previous sample-based methods Lin et al. (2022); Ye et al. (2022); Jang et al. (2023); Paranjape et al. (2023) that employ both instructions and instances. Moreover, instruction information alone shows a high correlation with task similarity as determined by complex pairwise transfer methods Zhou et al. (2023) and even achieves slightly better average zero-shot performance. This indicates that the instruction-based approach for selecting related tasks is not only easily applicable but also highly effective.
@@
We introduce an instruction-based task selection method for optimized instruction tuning. This method efficiently identifies relevant tasks without the extensive computation and data construction as in previous approaches.
By aligning InsTa with the instruction styles and formats of meta-datasets, we significantly improve performance, demonstrating the importance of understanding instructional nuances in instruction-based task selection.
Extensive experiments and comprehensive analyses across various benchmarks validate the superiority of our methodology, showcasing its enhanced efficiency, effectiveness, and practical applicability compared to previous methods.
@@ 2.1 Instruction Tuning: Generalist or Specialist
According to Kung and Peng (2023), instruction tuning as a generalist can be categorized based on its objectives: generalizing to "unseen tasks" and generalizing to "unseen instructions". Early studies propose the approach of generalizing to unseen tasks, which involves training on various natural language processing (NLP) tasks and evaluating on unseen tasks Wei et al. (2022); Sanh et al. (2022); Wang et al. (2022). However, the task-based approach to instruction tuning faces limitations as it is challenging to define all instructions as corresponding tasks, making it difficult to generalize to diverse, user-oriented instructions. In response, a methodology that trains LLMs on diverse instructions without clear task boundaries has been proposed, aiming to generalize to unseen instructions rather than unseen tasks Ouyang et al. (2022); Taori et al. (2023); Chiang et al. (2023).
@@ 2.2 Quantifying Task Relationship
The research on identifying task relationships has been extended to the field of instruction tuning as well. Zhou et al. (2023); Kim et al. (2023a) measure pairwise task transfer between every task pair to identify helpful source tasks on specific target tasks. On the other hand, Jang et al. (2023); Paranjape et al. (2023) have calculated task similarity using only a few data samples from training and evaluation tasks instead of training and evaluating every task pair individually.
Previous task selection approaches, which automatically measure task similarities, are time-consuming because they require training and evaluating every task pair or still rely on the availability of data samples. Our method simplifies this process by evaluating task relationships solely through instructions, eliminating the need for laborious measurements across every task or generating test data. For more detailed descriptions of our method’s practicalities and efficiency compared to prior works, please refer to Appendices A and B.
@@ 3.1 Formulation
where denotes the instruction of the target task , and denotes the instruction of some arbitrary task chosen for similarity assessment. For measuring similarity, we employ cosine similarity, and for the embedding function , we utilize the Sentence Transformer Reimers and Gurevych (2019) as an off-the-shelf embedding model, following Jang et al. (2023). For more specific details, please refer to the Appendix D.
@@ 3.2 Aligning Instruction-based Task Selector with Meta-dataset
The off-the-shelf embedding model often lacks the capability to accurately identify related tasks based on instructions, as it is not trained in the unique instruction styles present in meta-datasets. To mitigate this issue, our approach includes an additional aligning process that fine-tunes our selector model to adapt to the distinctive instruction styles of each meta-dataset. For training, we select a random instruction from the same task as the given instruction as a positive sample and designate instructions from different task clusters as negative samples.333The rationale behind not utilizing task instructions from the same task cluster as positive or negative samples is twofold: in the former case, even though tasks may have the same task type, they can still be clearly differentiated, and in the latter case, false negatives may occur, even within the same task type. The training objective is as follows:
@@ 4.1 Dataset
We conduct experiments on two representative instruction tuning meta-datasets: P3 (Public Pool of Prompts) Sanh et al. (2022); Bach et al. (2022) and NIV2 (SuperNaturalInstructions V2) Wang et al. (2022). P3 is a meta-dataset comprised of 12 task clusters. It contains a total of 35 tasks across 8 task clusters for training and 11 tasks across 4 task clusters for held-out evaluation. Note that each task in P3 includes 11.7 instructions on average. Conversely, NIV2 encompasses 72 task clusters of English tasks. The training set comprises a total of 756 tasks across 60 clusters, while the held-out evaluation tasks include 119 tasks across 12 task clusters. In the case of NIV2, each task consists of a single instruction. We additionally evaluate BIG-Bench bench authors (2023) and BIG-Bench Hard (BBH) Suzgun et al. (2023) as supplementary evaluation datasets. For more detailed information, please refer to Table 1 and Appendix F.1.
@@ 4.2 Task Selector Setup
instructions feature human-crafted, human-readable Task Definition, Positive Task Examples, Negative Task Examples, and Explanation. We utilize the Task Definition for training the task selector. Moreover, since NIV2 has only one instruction per task, there are no positive samples for training the task selector. To address this, we generate instructions for all NIV2 tasks using GPT-4 OpenAI (2023). Specifically, we generate a paraphrased one and employ it as a positive sample for training the task selector. For more information about a query used for GPT-4 API calls and generated instructions, please refer to the Appendix G.1.
@@ 4.3 Training Setup
Following previous studies Sanh et al. (2022); Wang et al. (2022), we employ all tasks, excluding those held out for evaluation, as our training set. In contrast to traditional approaches that train on all tasks, our strategy specifically trains only the top- tasks considered the most informative for each target task.
@@
We use the T5 LM-adapted model(3B) Lester et al. (2021) as our base model and train for 3 epochs with a constant learning rate of 1e-4 and 5e-5 for P3 and NIV2, respectively. We configure validation from training datasets and select the model that shows the best performance in the validation. Our experiment is in a true zero-shot setting, where no samples from the held-out task are used for checkpoint selection. Please refer to the Appendix E for more detail.
@@
For the P3 dataset, we use the following baselines. GPT-3(175B) is an autoregressive LM that has shown remarkable ability in following demonstrations provided in its instructions Brown et al. (2020). T0-11B / T0-3B have same architecture as T5 but trained on millions of samples from 35 different P3 tasks Sanh et al. (2022). T5(3B) is text-to-text pretrained LM without instruction tuning Raffel et al. (2020). T5(3B) + Random is trained on 5 random tasks from P3. T5(3B) + Pairwise Transfer is trained on top-5 tasks with the highest transferability scores demonstrated in Zhou et al. (2023), representing pairwise transfer task selection. T5(3B) + PE w/RoE denotes a model that selects relevant tasks by measuring cosine similarity between training and evaluation data samples Jang et al. (2023).
@@ 4.4 Evaluation Setup
For P3 evaluation, following the evaluation method from Sanh et al. (2022), we apply rank classification and measure the model’s performance on every instruction of the target task. We then calculate the average performance for the task. Note that each target task in P3 has 10.09 instructions on average. For NIV2 evaluation, we follow the same evaluation protocol as in Wang et al. (2022) and report ROUGE-L Lin (2004) score. We adopt greedy decoding with a maximum generation length of 256.
@@ 5.1 P3 Results
Table 2 shows experimental results on the 11 unseen datasets from P3. Compared to T0-3B, which is fully instruction-tuned on all 35 tasks from P3, the T5(3B) + Random model shows inferior performance, but both T5(3B) + InsTa model and T5(3B) + InsTa model outperform 10 tasks out of 11 tasks, with each exhibiting a performance gap of 4.83% and 7.10% on average, respectively. They also show superior performance compared to T5(3B) + PE w/ RoE, highlighting the instruction itself is sufficient enough to choose informative tasks without the use of data samples. Moreover, T5(3B) + InsTa outperforms T5(3B) + InsTa across 10 tasks, suggesting further aligning task selector on instructions from the meta-dataset increases the precision of task selection. Additionally, our approach exhibits marginally superior performance compared to T5(3B) + Pairwise Transfer, suggesting that instruction-only task selection can effectively identify related tasks without the need for exhaustive pairwise evaluations. Notably, we observe a strong correlation between pairwise task transferability and our task selection approach, further elaborated in Appendix H.
@@ 5.2 Big-Bench Results
Table 3 evaluates the performance of instruction-based task selection on the 14 Big-Bench tasks employing P3 held-in tasks in Section 5.1 as the training dataset. We compare our method to baselines T0-3B and T5(3B) + Cos PE of Jang et al. (2023). Our model, T5(3B) + InsTa, surpasses T0-3B in 12 out of 14 tasks, achieving an average performance increase of 5.94%. Tasks like Movie D. and Syllogism, where results are comparable to random guesses, indicate informative tasks are insufficient in P3. When compared to Cos PE Jang et al. (2023), our T5(3B) + InsTa model shows improvements in most tasks, with an average increase of 2.07%. Notably, it also surpasses T0-11B in a majority of tasks. These findings demonstrate that instruction-based task selection enables more effective training even with a small number of tasks.